OpenStudy (lamp54321):

A meteor streaking through the night sky is located with radar. At point A its coordinates are (5.60 km , 1.35 km ), and 1.16 s later its has moved to point B with coordinates (6.08 km , 0.920 km ). 1) Find the x component of its average velocity between A and B 2) Find the y component of its average velocity between A and B. 3) Find the magnitude of its average velocity between these two points. 4) Find the direction of its average velocity between these two points.

OpenStudy (sapphiremoon):

A nice vector question before bed. :) The x component of the average velocity can be found by subtracting the x coordinate of A from that of B, 6.08 - 5.60 = 0.48. The same can be done for the y. 0.920 - 1.35 = -0.43. The magnitude of the average velocity is the magnitude of the vector with those x and y components: \[\sqrt{0.48 ^{2}+(-0.43) ^{2}} = 0.644\] The direction of the average velocity is the angle of this vector, notably: \[\theta = \tan^{-1} (\frac{-0.43}{ 0.48 }) = -41.9\] Do you need any further explanation?

OpenStudy (raffle_snaffle):

^Next time you should let OP answer the questions.=)

OpenStudy (sapphiremoon):

@raffle_snaffle Well, I was trying to explain the answer... Apparently people like it better when you bridge the entire gap instead of leaving the last step open? Idk, just trying to do my best here.